Anguilla Manufacturing is interested in using the data collected during value stream mapping to evaluate the current state of performance of the capacity of its manual assembly line process under different batch size assumptions. The availability of processing per shift, after accounting for breaks and a midshift employee meeting, is 436 minutes. The current operating characteristics of each processing step are found in the accompanying table. Note that each step can only process one part at a time and all steps must process the same sized batches.

Cycle time per part: 20 seconds (saw), 15 seconds (sand), 30 seconds (drill), 25 seconds (assemble), 10 seconds (mark)

setup time per batch: 3 minutes (saw), 4 minutes (sand), 0 minutes (drill), 3 minutes (assemble), 8 minutes (mark)

a) Calculate the average processing time per unit and the capacity at each step assuming batch sizes of: 10 units, 20 units, 30 units, 40 units.

b) Identify the bottleneck oepration and the line's processing capacity for each batch size listed in part (a)

c) explain why batch sizes beyond 40 units will not increase the line's processing capacity further.  

Thanks in advance!
Answer #1

a) For the given batch sizes, average processing time per unit is computed at each step as below:

D13 a fx LOOKUP(D10,D$3:D$7,$A$3:$A$7) Average processing time per unit for batch size, a 2 Step 3 Saw 4 Sand 5 Drill 6 Assemble 7 Mark Cycle Setup time 180 240 10 30 26.0 23.0 30.0 31.0 26.0 20 time 20 15 30 25 10 38.0 39.0 30.0 43.0 58.0 29.0 27.0 30.0 34.0 34.0 24.5 21.0 30.0 29.5 22.0 180 480 10 Longest processing time- 31 30 12 13 14 15 16 17 18 19 20 Bottleneck step Mark Mark Assemble Drill

EXCEL Formulas:

b) Bottleneck operation is identified in part (a) above.

Processing capacity = 1 / Cycle time, where Cycle time is the average processing time of bottleneck step.

For batch size of 10, processing capacity = (1/58)*60 = 1.034 per minute

For batch size of 20, processing capacity = (1/34)*60 = 1.765 per minute

For batch size of 30, processing capacity = (1/31)*60 = 1.935 per minute

For batch size of 40, processing capacity = (1/30)*60 = 2.000 per minute

c) The average processing time of Drill is constant at 30 sec irrespective of batch size, because it does not have any setup time.

With a batch size of 40 units, the cycle time of 30 second is achieved, which is equal to the average processing time of Drill. So, Drill is the bottleneck step for a batch size of 40 units. So further, increasing the batch size, will not reduce the cycle time of Drill operation, as already elaborated.

Hence, the line's processing capacity will not improve further, despite increasing the batch size beyond 40 units.